# A Fast and Model Based Approach for Evaluating Task-Competence of Antagonistic Continuum Arms

### v1

@@ I INTRODUCTION
Figure 1: We propose a novel method to evaluate fluid-driven soft arm designs on their ability to meet a task shape at the task load (a). In this work, we show the method for planar antagonistic architectures (b). Each segment of the arm has two extending and two contracting actuators, and arms move via selective actuator pressurization.
Figure 2: Overview of our mechanics model. A soft robot arm composed of two bellows and two muscles can be fully parameterized by its centerline twists $\accentset{\scriptscriptstyle{\circleright}}{g}_{oi}$ and transformations between actuators ${}_{o}g_{o\alpha}$ (a). When a tip load $\mathbf{q}_{\textrm{tip}}$ is applied, wrenches $\mathbf{q}_{i}$ are induced along the arm’s backbone. To achieve static equilibrium, each actuator contributes a reaction force and moment to balance the load (b). We consider two types of actuators: contracting McKibben artificial muscles and extensile bellows actuators. The characterized force functions $f(\epsilon,p)$ of each actuator are also shown, with actuation regimes labeled according to [17] (c).
@@ II CONTINUUM MODEL
    Bending Stiffness: We assume the arm has bending stiffness beyond that generated by the uniaxial actuator forces, and is dependent on curvature and pressure.
@@ II-A Cosserat rod kinematics
Consider a robot arm with $M$ actuators $\{A,B,\dots,\mu\}$ discretized by $N+1$ nodes into $N$ constant curvature segments ($\mu$ stands for the $M$-th actuator). Note that discretization is for numerical solution, and does not represent physical segmentation. Each actuator $\alpha$ is mounted at a distance of $r_{\alpha}$ away from the center. For any actuator $\alpha$, the pose of its $i$-th node can be denoted as ${}_{\alpha}g_{i}=\begin{bmatrix}{}_{\alpha}x_{i}&{}_{\alpha}y_{i}&{}_{\alpha}\theta_{i}\end{bmatrix}^{\textrm{T}}$, with matrix representation:
@@
We will notate pose composition as $g_{B}=g_{A}\circ g_{AB}$, which is implemented as the product of the matrix forms $\rho(g_{B})=\rho(g_{A})\rho(g_{AB})$. At each node $i$ of actuator $\alpha$, the twist vector ${}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{i}=\begin{bmatrix}{}_{\alpha}l_{i}&{}_{\alpha}\gamma_{i}&{}_{\alpha}\kappa_{i}\end{bmatrix}$ describes the body-frame rate of change of $g_{\alpha i}$ with $l$ as the instantaneous length of the segment, $\gamma$ the shear, and $\kappa$ the curvature. The matrix representation of the twist vector is:
@@
These twists can be integrated using the exponential map $\exp(\accentset{\scriptscriptstyle{\circleright}}{g})=\exp_{M}(\rho(\accentset{\scriptscriptstyle{\circleright}}{g}))$ to recover poses along each actuator:
${}_{\alpha}g_{i}={}_{\alpha}g_{0}\circ\prod_{k=1}^{i}\exp(\frac{1}{n}\accentset{\scriptscriptstyle{\circleright}}{{}_{\alpha}g_{k}})$ (3)
Using eqn. 3, we can describe the poses along each actuator ${}_{\alpha}g=\{{}_{\alpha}g_{1},{}_{\alpha}g_{2},\cdots{}_{\alpha}g_{N+1}\}$ using the base pose ${}_{\alpha}g_{1}$ and the twists of each segment ${}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}=\{{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{1},{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{2},\cdots,{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{N}\}$, which are more closely related to the mechanics (Fig. 2A). Furthermore, our assumption of a rigid and constant cross-section enables us relate the twists of each actuator to the twists of the manipulator centerline. Let the pose and twists of the manipulator centerline at node $i$ be ${}_{o}g_{i}$ and ${}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}_{i}$ respectively, and let ${}_{o}g_{o\alpha}$ be the transformation from actuator $o$ to actuator $\alpha$ in each cross-section, which is usually $\begin{bmatrix}0&r_{\alpha}&0\end{bmatrix}^{\textrm{T}}$. As shown in [19], the twist of any actuator can be computed from the centerline twist as:
@@ II-B External Loads
Static equilibrium is achieved by balancing internal reactions to external loading. We will refer to the combination of a force and moment at a point as a wrench, and consider external wrenches applied to the tip of a robot arm $\mathbf{q}_{\textrm{tip}}=\begin{bmatrix}f_{x}&f_{y}&m\end{bmatrix}$. This tip wrench will induce a wrench at every node along the arm’s centerline with equal force but additional moment due to the moment arm (see Fig. 3A-left). The wrench at each node $i$ along the arm’s centerline is:
@@ III-A Problem Statement
Given an arm design as defined above, we are interested in its ability to sustain desired shapes when carrying specific payloads. Specifically, when an arm is subject to a tip load $\mathbf{q}_{\textrm{tip}}^{*}$, we ask whether it can maintain a desired equilibrium centerline shape ${}_{o}g^{*}=\{{}_{o}g^{*}_{1},{}_{o}g^{*}_{2},\dots,{}_{o}g^{*}_{N}\}$, or equivalently twists ${}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*}=\{{}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*}_{1},{}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*}_{2},\dots,{}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*}_{N}\}$. We define the pair $(\mathbf{q}_{\textrm{tip}}^{*},{}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*})$ as a task, and refer to its components respectively as the task load and task shape.
An arm’s control space is determined by the pressure limits of each actuator. If each actuator has max pressure $\bar{p}_{A},\bar{p}_{B},\dots,\bar{p}_{\mu}$ and minimum pressure zero, then the input pressure space is $\mathcal{P}=[0,\bar{p}_{A}]\times[0,\bar{p}_{B}]\times\dots\times[0,\bar{p}_{\mu}]$. When a proposed arm design is capable of completing a desired task, this means that there exists a pressure $\mathbf{p}^{*}$ in $\mathcal{P}$ such that the arm’s equilibrium shape when subject to the task load is equal to the specified task shape, i.e. ${}_{o}\hat{g}(\mathbf{p}^{*},\mathbf{q}_{\textrm{tip}}^{*})={}_{o}g^{*}$. If this is true, we say the task is attainable.
@@ III-C1 Absolute Attainability
Our first condition is to check that each absolute requirement wrench $\mathbf{w}_{i}^{*}$ lies within each $\mathbf{a}_{i}[\mathcal{P}]$, which we call the "$i$-th attainable wrench space". We claim the following lemmas to aid computing each $\mathbf{a}_{i}[\mathcal{P}]$:
@@
These lemmas are consequences of our actuator models being continuous and monotonic, which preserve topological invariants like boundaries and connectedness [20] (see SI for proof). We can therefore approximate whether $\mathbf{w}^{*}_{i}$ is in $\mathbf{a}_{i}[\mathcal{P}]$ by sampling a set of points $\mathbf{E}$ from the edges of $\partial\mathcal{P}$ and checking if $\mathbf{w}^{*}_{i}$ is enclosed by the convex hull $\textrm{Conv}(\mathbf{a}_{i}[{\mathbf{E}}])$, which we call $\mathcal{H}_{i}$, the $i$-th absolute wrench hull, which also visualizes the possible reactions across $\mathcal{P}$. Across all $i$, we will refer to the collection $\mathcal{H}=(\mathcal{H}_{1},\dots,\mathcal{H}_{N})$ as the absolute wrench hull, and if each $\mathbf{w}_{i}^{*}$ is enclosed in each $\mathcal{H}_{i}$ then we say the task $({}_{o}\accentset{\scriptscriptstyle{\circleright}}{g},\mathbf{q}_{\textrm{tip}}^{*})$ is absolutely attainable (Fig. 3D). If any $\mathbf{w}^{*}_{i}$ are not in $\textrm{Conv}(\mathbf{a}_{i}[\mathbf{E}])$, the shortest distance between them can be found by solving a convex quadratic program described in our SI, and we will refer to their sum as the absolute unattainability.
@@ III-C2 Relative Attainability
Absolute attainability is insufficient alone to guarantee that a task is truly attainable. If a requirement wrench sequence does not change along $i$ in the same "direction" as the attainable wrench sequences - as in Fig. 3D - then it cannot be truly attainable. We now formalize this "direction" as a second test for attainability.
Consider the difference $\Delta\mathbf{w}_{i}^{*}$ of each $\mathbf{w}_{i}^{*}$ from its starting point $\mathbf{w}_{1}^{*}$. We will refer to the sequence of such differences i.e., $\Delta\mathbf{w}^{*}=(\mathbf{0},\mathbf{w}_{2}^{*}-\mathbf{w}_{1}^{*},\ldots,\mathbf{w}_{N}^{*}-\mathbf{w}_{1}^{*})$, as the relative requirement wrench sequence. For any attainable wrench sequence $\mathbf{a}(\mathbf{p})$, a similar sequence $\Delta\mathbf{a}(\mathbf{p})$ can also be defined, and we call it the relative attainable wrench sequence. The collection of such sequences over $\mathcal{P}$ is the space of relative attainable wrench sequences, and is defined as:
@@
For a task to be attainable, there must exist $\mathbf{p}$ so that $\mathbf{a}(\mathbf{p})=\mathbf{w}^{*}$, which implies that $\Delta\mathbf{a}(\mathbf{p})=\Delta\mathbf{w}^{*}$. Determining if such a $\Delta\mathbf{a}(\mathbf{p})$ exists is equivalent to checking if $\Delta\mathbf{w}^{*}$ lies within $\Delta\mathbf{a}[\mathcal{P}]$, which is difficult. We will again approximate the problem by individually considering if each $\Delta\mathbf{w}_{i}^{*}$ lies within its respective $\Delta\mathbf{a}_{i}[\mathcal{P}]=\{\mathbf{a}_{i}(\mathbf{p})-\mathbf{a}_{1}(\mathbf{p}\mid\mathbf{p}\in\mathcal{P}\}$. As with the absolute attainability, we note the following properties of each $\Delta\mathbf{a}_{i}[\mathcal{P}]$:
@@
A formal proof of these properties eludes us, but an extensive search yielded no counterexamples (see demonstrations in Fig. 3E and 4). As before, we can approximate whether $\Delta\mathbf{w}_{i}^{*}$ is in $\Delta\mathbf{a}_{i}[\mathcal{P}]$ by checking if $\Delta\mathbf{w}_{i}^{*}$ is enclosed by $\textrm{Conv}(\Delta\mathbf{a}_{i}[\mathbf{E}])$, which we name the $i$-th relative wrench hull. If true for all $i$, then the task is relatively attainable. If any $\Delta\mathbf{w}_{i}^{*}$ are not in $\textrm{Conv}(\Delta\mathbf{a}_{i}[\mathbf{E}])$, the shortest distance between them can be found by solving a convex quadratic program, and we will refer to their sum as the relative unattainability.
@@ IV-B Validation of Wrench-Hull Analysis
To confirm that tasks assigned a low absolute unattainability for a given arm are indeed more feasible for that arm, for each task-load we used the search-based method described in section III-B to find each arm’s best attempt to match the task-shape while subject to each of the 67 task-loads. The resulting equilibrium arm shapes are shown in Fig. 5D and E, with each shape’s correspondent task-loads shown in yellow. As predicted by the absolute unattainability metric, the antagonistic arm’s equilibrium shapes have less variance than the bellows-only arm thanks to its ability to increase stiffness to withstand larger loads. Furthermore, by highlighting each arm’s equilibrium shapes for task-loads that have below-median unattainability, we see that without exception the tasks with lower unattainability are indeed closer to being accomplished.
@@ IV-C Evaluation of Antagonistic Actuation
The antagonistic arm has the lowest absolute unattainability for the first two task-shapes, followed by the bellows-only arm, and finally the McKibben-only arm. Comparing the absolute wrench-hulls of each arm offers insights for why the antagonistic arm is more capable of the two: because the antagonistic arm combines extending and contracting actuators, its absolute wrench-hull is centered near the origin, and therefore more capable of enclosing requirement wrench sequences, which are also centered around the origin. The bellows-only arm may have a larger wrench-hull, but because it only has extending actuators its wrench-hull is only positioned in the positive reaction-force half-plane. Likewise the McKibben-only arm, whose actuators can only contract, has a wrench-hull entirely with negative reaction-force. Passive actuator reaction forces do provide countering force, as noted in [18], but are strain-dependent and small in magnitude, while antagonistic countering forces are controllable via pressure and can be large in magnitude.
On the third shape, the bellows-only arm has the lowest absolute unattainability, with the antagonistic arm a distant second. This is because the large curvature near the tip of the arm is difficult to attain with an antagonistic arm, which has higher stiffness. The bellows-only arm has lower stiffness (note the drooping for high downward loads in Fig. 5D) and thus can better match the shape.

### v3

@@ I INTRODUCTION
Figure 1: (A): We propose a method to evaluate fluid-driven soft arm designs on their ability to sustain a task shape at the task load. (B): We mainly demonstrate our method with planar antagonistic arm designs. Each arm segment has two extending and two contracting actuators, and arms move via selective actuator pressurization.
Figure 2: Overview of our mechanics model. (A): A soft robot arm composed of two bellows and two muscles can be fully parameterized by its centerline twists $\accentset{\scriptscriptstyle{\circleright}}{g}_{oi}$ and transformations between actuators ${}_{o}g_{o\alpha}$. (B): When a tip load $\mathbf{q}_{\textrm{tip}}$ is applied, wrenches $\mathbf{q}_{i}$ are induced along the arm’s backbone. To achieve static equilibrium, each actuator contributes a reaction force and moment to balance the load. (C): We consider two types of actuators: contracting McKibben artificial muscles and extensile bellows actuators. The characterized force functions $f(\epsilon,p)$ of each actuator are also shown, with actuation regimes labeled according to [17].
@@ II CONTINUUM MODEL
    Bending Stiffness: We assume the actuators have bending stiffness beyond that generated their uniaxial forces, and is dependent on curvature and pressure.
@@ II-A Cosserat rod kinematics
Consider a robot arm with $M$ actuators $\{A,B,\dots,\mu\}$ discretized by $N+1$ nodes into $N$ constant curvature segments ($\mu$ stands for the $M$-th actuator). Note that discretization is for numerical solution, and does not represent physical segmentation. Let $\alpha$ be any actuator. Each actuator $\alpha$ is mounted at a distance of $r_{\alpha}$ away from the center. For any actuator $\alpha$, the pose of its $i$-th node can be denoted as ${}_{\alpha}g_{i}=\begin{bmatrix}{}_{\alpha}x_{i}&{}_{\alpha}y_{i}&{}_{\alpha}\theta_{i}\end{bmatrix}^{\textrm{T}}$, with matrix representation:
@@
We will notate pose composition as ${}_{\alpha}g_{2}={}_{\alpha}g_{1}\circ{}_{\alpha}g_{12}$, which is implemented as the product of the matrix forms $\rho({}_{\alpha}g_{2})=\rho({}_{\alpha}g_{1})\rho({}_{\alpha}g_{12})$. At each node $i$ of actuator $\alpha$, the twist vector ${}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{i}=\begin{bmatrix}{}_{\alpha}l_{i}&{}_{\alpha}\gamma_{i}&{}_{\alpha}\kappa_{i}\end{bmatrix}^{\textrm{T}}$ describes the body-frame rate of change of $g_{\alpha i}$ with $l$ as the instantaneous length of the segment, $\gamma$ the shear, and $\kappa$ the curvature. The matrix representation of the twist vector is:
@@
These twists can be integrated using the exponential map $\exp(\accentset{\scriptscriptstyle{\circleright}}{g})=\exp_{M}(\rho(\accentset{\scriptscriptstyle{\circleright}}{g}))$ to recover poses along each actuator, where $\exp_{M}$ is the matrix exponential:
${}_{\alpha}g_{i}={}_{\alpha}g_{1}\circ\prod_{k=1}^{i-1}\exp(\frac{1}{n}\accentset{\scriptscriptstyle{\circleright}}{{}_{\alpha}g_{k}})$ (3)
Using eqn. 3, we can describe the poses along each actuator ${}_{\alpha}g=\{{}_{\alpha}g_{1},{}_{\alpha}g_{2},\cdots{}_{\alpha}g_{N+1}\}$ using the base pose ${}_{\alpha}g_{1}$ and the twists of each segment ${}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}=\{{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{1},{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{2},\cdots,{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{N}\}$, which are more closely related to the mechanics (Fig. 2A). Furthermore, our assumption of a rigid and constant cross-section enables us relate the twists of each actuator to the twists of the manipulator centerline. Let the pose and twists of the manipulator centerline at node $i$ be ${}_{o}g_{i}$ and ${}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}_{i}$ respectively, and let ${}_{o}g_{o\alpha}$ be the transformation from centerline $o$ to actuator $\alpha$ in each cross-section, which is usually $\begin{bmatrix}0&r_{\alpha}&0\end{bmatrix}^{\textrm{T}}$. As shown in [19], the twist of any actuator can be computed from the centerline twist as:
@@ II-B External Loads
Static equilibrium is achieved by balancing internal reactions to external loading. We will refer to the combination of a force and moment at a point as a wrench, and consider external wrenches applied to the tip of a robot arm $\mathbf{q}_{\textrm{tip}}=\begin{bmatrix}f_{x}&f_{y}&m\end{bmatrix}^{\textrm{T}}$. This tip wrench will induce a wrench at every node along the arm’s centerline with equal force but additional moment due to the moment arm (see Fig. 3A-left). The wrench at each node $i$ along the arm’s centerline is:
@@ III-A Problem Statement
Given an arm design as defined above, we are interested in its ability to sustain desired shapes when carrying specific payloads. Specifically, when an arm is subject to a tip load $\mathbf{q}_{\textrm{tip}}^{*}$, we ask whether it can maintain a desired equilibrium centerline shape ${}_{o}g^{*}=\{{}_{o}g^{*}_{1},{}_{o}g^{*}_{2},\dots,{}_{o}g^{*}_{N+1}\}$, or equivalently twists ${}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*}=\{{}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*}_{1},{}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*}_{2},\dots,{}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*}_{N}\}$. We define the pair $(\mathbf{q}_{\textrm{tip}}^{*},{}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}^{*})$ as a task, and refer to its components respectively as the task load and task shape.
An arm’s control space is defined by the pressure limits of each actuator. If each actuator has max pressure $\bar{p}_{A},\bar{p}_{B},\dots,\bar{p}_{\mu}$ and minimum pressure zero, then the input pressure space is $\mathcal{P}=[0,\bar{p}_{A}]\times[0,\bar{p}_{B}]\times\dots\times[0,\bar{p}_{\mu}]$. When a proposed arm design is capable of completing a desired task, this means that there exists a pressure $\mathbf{p}^{*}$ in $\mathcal{P}$ such that the arm’s equilibrium shape when subject to the task load is equal to the specified task shape, i.e. ${}_{o}\hat{g}(\mathbf{p}^{*},\mathbf{q}_{\textrm{tip}}^{*})={}_{o}g^{*}$. If this is true, we say the task is attainable.
@@ III-C1 Absolute Attainability
Our first condition is to check that each absolute requirement wrench $\mathbf{w}_{i}^{*}$ lies within each $\mathbf{a}_{i}[\mathcal{P}]$, which we call the "$i$-th attainable wrench space". We claim the following propositions for computing each $\mathbf{a}_{i}[\mathcal{P}]$:
@@
These propositions are consequences of our actuator models being continuous and monotonic, which preserve topological invariants like boundaries and connectedness [20] (see SI for proof). We can therefore approximate whether $\mathbf{w}^{*}_{i}$ is in $\mathbf{a}_{i}[\mathcal{P}]$ by sampling a set of points $\mathbf{E}$ from the edges of $\partial\mathcal{P}$ and checking if $\mathbf{w}^{*}_{i}$ is enclosed by the convex hull $\textrm{Conv}(\mathbf{a}_{i}[{\mathbf{E}}])$. We call this $\mathcal{H}_{i}$, the $\mathit{\mathbf{i}}$-th absolute wrench hull, which also visualizes the possible reactions across $\mathcal{P}$. Across all $i$, we will refer to the collection $\mathcal{H}=(\mathcal{H}_{1},\dots,\mathcal{H}_{N})$ as the absolute wrench hull, and if each $\mathbf{w}_{i}^{*}$ is enclosed in each $\mathcal{H}_{i}$ then we say the task $({}_{o}\accentset{\scriptscriptstyle{\circleright}}{g},\mathbf{q}_{\textrm{tip}}^{*})$ is absolutely attainable (Fig. 3D). If any $\mathbf{w}^{*}_{i}$ are not in $\mathcal{H}_{i}$, the shortest distance between them can be found by solving a convex quadratic program described in our SI, and we will refer to their sum as the absolute unattainability.
@@ III-C2 Relative Attainability
Absolute attainability is insufficient alone to guarantee that a task is truly attainable. If a requirement wrench sequence does not change along $i$ in the same "direction" as any attainable wrench sequences - as in Fig. 3D - then it cannot be truly attainable. We now formalize this "direction" as a second test for attainability.
Consider the difference $\Delta\mathbf{w}_{i}^{*}$ of each $\mathbf{w}_{i}^{*}$ from its starting point $\mathbf{w}_{1}^{*}$. We will refer to the sequence of such differences i.e., $\Delta\mathbf{w}^{*}=(\mathbf{0},\mathbf{w}_{2}^{*}-\mathbf{w}_{1}^{*},\ldots,\mathbf{w}_{N}^{*}-\mathbf{w}_{1}^{*})$, as a relative requirement wrench sequence. For any attainable wrench sequence $\mathbf{a}(\mathbf{p})$, a similar sequence $\Delta\mathbf{a}(\mathbf{p})$ can also be defined, and we call it the relative attainable wrench sequence. The collection of such sequences over $\mathcal{P}$ is the space of relative attainable wrench sequences, and is defined as:
@@
For a task to be attainable, there must exist $\mathbf{p}$ so that $\mathbf{a}(\mathbf{p})=\mathbf{w}^{*}$, which implies that $\Delta\mathbf{a}(\mathbf{p})=\Delta\mathbf{w}^{*}$. Determining if such a $\Delta\mathbf{a}(\mathbf{p})$ exists is equivalent to checking if $\Delta\mathbf{w}^{*}$ lies within $\Delta\mathbf{a}[\mathcal{P}]$, which is difficult. We will again approximate the problem by individually considering if each $\Delta\mathbf{w}_{i}^{*}$ lies within its respective $\Delta\mathbf{a}_{i}[\mathcal{P}]=\{\mathbf{a}_{i}(\mathbf{p})-\mathbf{a}_{1}(\mathbf{p}\mid\mathbf{p}\in\mathcal{P}\}$. As with the absolute attainability, we claim the following properties of each $\Delta\mathbf{a}_{i}[\mathcal{P}]$:
@@
An extensive search yielded no counterexamples to these two properties, and their proof will be provided in a future work (see demonstrations in Fig. 3E and 4). As before, we can approximate whether $\Delta\mathbf{w}_{i}^{*}$ is in $\Delta\mathbf{a}_{i}[\mathcal{P}]$ by checking if $\Delta\mathbf{w}_{i}^{*}$ is enclosed by $\textrm{Conv}(\Delta\mathbf{a}_{i}[\mathbf{E}])$, which we name the $i$-th relative wrench hull. If true for all $i$, then the task is relatively attainable. If any $\Delta\mathbf{w}_{i}^{*}$ are not in $\textrm{Conv}(\Delta\mathbf{a}_{i}[\mathbf{E}])$, the shortest distance between them can be found by solving a convex quadratic program, and we will refer to their sum as the relative unattainability.
@@ IV-B Validation of Wrench-Hull Analysis
To confirm that tasks assigned a low absolute unattainability for a given arm are indeed more feasible for that arm, we used the search-based method described in section III-B to find each arm’s best attempt to match the task-shape while subject to each of the 67 task-loads. The resulting equilibrium arm shapes are shown in Fig. 5D and E, with each shape’s correspondent task-loads shown in yellow. As predicted by the absolute unattainability metric, the antagonistic arm’s equilibrium shapes have less variance than the bellows-only arm thanks to its ability to increase stiffness to withstand larger loads. Furthermore, highlighting each arm’s equilibrium shapes for task-loads that have below-median unattainability shows that, without exception, the tasks with lower unattainability are closer to being accomplished.
@@ IV-C Evaluation of Antagonistic Actuation
The antagonistic arm has the lowest absolute unattainability for the first two task-shapes, followed by the bellows-only arm, and finally the McKibben-only arm. Comparing the absolute wrench-hulls of each arm suggests why the antagonistic arm is more capable of the two: because the antagonistic arm combines extending and contracting actuators, its absolute wrench-hull is closer to the origin, and therefore better at enclosing requirement wrench sequences, which are centered around the origin. The bellows-only arm may have a larger wrench-hull, but because it only has extending actuators, its wrench-hull only extends to the positive reaction force half-plane. Likewise the McKibben-only arm, whose actuators can only contract, has a wrench-hull entirely with negative reaction-force. Passive actuator reaction forces do provide countering force, as noted in [18], but are strain-dependent and small in magnitude, while antagonistic countering forces are controllable via pressure and can be large in magnitude.
On the third shape, the bellows-only arm has the lowest absolute unattainability, with the antagonistic arm a distant second. This is because the large curvature near the tip of the arm is difficult to attain with an antagonistic arm, which has higher stiffness. The bellows-only arm has lower stiffness (see Fig. 5D) and thus can better match the shape.